OpenStudy (anonymous):

20 minus (-14) = A. -34 B. 34 C. -6 D. 6

OpenStudy (anonymous):

20-(-14) both negative will become +ive

OpenStudy (anonymous):

20-(-14)=20+14=36 B

OpenStudy (anonymous):

34 sorry